Fermi National Laboratory
 

Accelerator Update for July 6, 2007 - July 9, 2007

 

Friday July 6

The day shift began with Operations monitoring the Tevatron (TeV) store 5540 (80.8E30) and the Recycle stash (218.1E10), with the Antiproton Source (Pbar) stacking (40.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:14 AM, two MI abort kickers (MIKPS4A & B) tripped off for the first of two times in nine minutes. Experts investigated.

MI experts tried to begin a study period at 11:14 AM, but communication errors were hindering their efforts. An expert investigated.

At 11:50 AM, per instruction, Operations rebooted MECAR to resolve the MI communication errors.

At 1:42 PM, Operations reported that the Pelletron had tripped off over 20 times. Operations contacted an expert.

Operations established beam to the new Meson MTest experiment (T964) at 4:01 PM.

Operations reported at 4:01 PM that the Recycler had stashed antiprotons three times during the day shift.

At 4:30 PM, Operations reported a power glitch. The RF went down in Linac, Booster, and MI and various power supplies tripped off. Fortunately the stack, stash, and store all survived.

Operations recovered from the power glitch and had beam restored by 4:43 PM.

Operations disabled MTest for the evening at 6:04 PM.

Operations terminated store 5540 at 7 PM.

Operations put the TeV into shot setup at 7:14 PM.

Operations established store 5541 at 9:11 PM with an initial luminosity of 205.79E30.

Operations reported at 11:50 PM that the Recycler had stashed three times during the evening shift.

Saturday July 7

The midnight shift began with Operations monitoring store 5541 (128.14E30) and the stash (90.1E10), with Pbar stacking (13.0E10), and with NuMI and MiniBooNE taking beam.

At 1:12 AM, the NuMI beam permit tripped off due to a MI RF problem. It reset.

At 1:13 AM, Operations reported that the stack disappeared. Operators investigated and discovered that the output of an Accumulator dipole shunt had spiked down at the same time the stack was lost. Operators informed experts.

At 2:56 AM, a Linac Buncher tripped off. It reset.

Operations established beam to MTest at 5:59 AM.

At 7:55 AM, BRF11 tripped off. It reset.

Operations reported at 7:55 AM that the Recycler had stashed twice during the midnight shift.

At 10:53 AM, Operations reported that the H- Source output current been decreasing. An operator made an adjustment.

At 1:29 PM, MIRF11 tripped off for the first of six times in four hours. Each time the RF station tripped, the NuMI beam permit tripped. Everything reset. Operations contacted an expert.

At 5:53 PM, the Tevatron Electron Lens (TEL2) tripped off. It reset.

At 6:09 PM, Operations reported that they were still having problems with MIRF11 and that MI beam was off while operators consulted with an expert. Operators made adjustments.

At 6:24 PM, MIRF11 again tripped off. Operators bypassed the station.

Operations disabled MTest at 7:09 PM.

At 7:46 PM, operators investigated an indication that the MI interlocks had dropped.

Operations reported at 8:01 PM that the Recycler had stashed five times during the last 12 hours. Accelerator status: store 5541 was at 37.8E30, the stack was at 19.4E10, and the stash was at 307.9E10.

At 8:10 PM, the Linac Buncher tripped off. It reset.

At 9:20 PM, LRF2 tripped off for the first of two times in 40 minutes. It reset.

Operations terminated store 5541 at 9:58 PM.

At 10:12 PM, the Linac Buncher tripped off for the first of two times in about an hour. It reset.

Operations put the TeV into shot setup at 10:21 PM.

At 11:42 PM, Operations reported a possible LCW leak in Pbar. Operators investigated.

Operations established store 5542 at 11:52 PM with an initial luminosity of 188.4E30.

Sunday July 8

The midnight shift began with Operations monitoring store 5542 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations prepared Pbar for access at 12:16 AM. Operators searched Pre-Vault, Pre-Target, and Transport, but found no leak. At 12:57 AM, operators reported finding a hose leak on a Pbar Debuncher quadrupole magnet (D4Q6). An expert had already been called in and he began repairs.

Operations reported at 1:23 AM that the D4Q6 hose leak was repaired.

Pbar resumed stacking at 2:35 AM.

Around 4 AM, CDF reported that TEL2 had tripped off. Operators investigated and found that it had tripped off on a vacuum interlock. Everything reset.

MTest resumed taking beam at 6:21 AM.

Operations reported at 8:18 AM that the Recycler had stashed twice during the last 12 hours. Accelerator status: store 5542 was at 77.52E30, the stack was at 44.6E10, and the stash was at 86.7E10.

At 1:09 PM, the Meson M02-5 electrical safety system permits dropped. Operations called in an interlock expert.

An interlock expert reported at 3:16 PM that he had replaced the Meson safety system’s power supply.

At 3:42 PM, LRF3 tripped off. Operators investigated and smelled burnt electronics. Operations called in an expert.

Operators reported that they had search and secured the M02-5 enclosures by 4:11 PM.

At 5:07 PM, the M02-5 interlocks dropped again. Operations called in experts.

Operations resumed sending beam to Pbar, NuMI, and MiniBooNE at 7:05 PM.

The Linac expert reported at 7:11 PM that he found and replaced a failed modulator resistor board and a crowbar switch relay.

At 7:23 PM, LRF3 tripped off again. The Linac expert was called back in.

Operations reported at 8 PM that the Recycler had stashed three times during the last 12 hours. Accelerator status: store 5542 was at 41.06E30, the stack was at 45.1E10, and the stash was at 180.8E10.

Interlock experts reported at 9:06 PM that they investigated and then replaced the power supply again.

With the completion of LRF3 repairs, Linac resumed sending beam at 11 PM.

Monday July 9

The midnight shift began with Operations monitoring store 5542 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:26 AM LRF3 tripped off. It reset.

At 12:47 AM, the MI-8 collimator tripped off. It reset.

Operations terminated store 5542 at 12:59 AM.

Operations put the TeV into shot setup at 1:13 AM.

Operations established store 5543 at 3:46 AM with an initial luminosity of 150.15E30.

Operators reported at 6:20 AM that they had searched and secured M02-5.

Operations established beam to MTest at 7 AM.

Operations reported at 8:11 AM that the Recycler had stashed twice during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow a few studies during the day. There will most likely be a CDF two-shift access beginning at 6 AM on Wednesday, 7/11/07, so CDF can make necessary repairs. The Run Coordinator will allow certain machine work during this access.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/09/2007    email Fermilab